### v1

# Flowception: Temporally Expansive Flow Matching for Video Generation

Abstract
We present Flowception, a novel non-autoregressive and variable-length video generation framework. Flowception learns a probability path that interleaves discrete frame insertions with continuous frame denoising. Compared to autoregressive methods, Flowception alleviates error accumulation/drift as the frame insertion mechanism during sampling serves as an efficient compression mechanism to handle long-term context. Compared to full-sequence flows, our method reduces FLOPs for training three-fold, while also being more amenable to local attention variants, and allowing to learn the length of videos jointly with their content. Quantitative experimental results show improved FVD and VBench metrics over autoregressive and full-sequence baselines, which is further validated with qualitative results. Finally, by learning to insert and denoise frames in a sequence, Flowception seamlessly integrates different tasks such as image-to-video generation and video interpolation.
[…]
Figure 1: Examples of image-to-video (I2V) generation and video interpolation with Flowception . Input frames marked by dashed boundaries. Flowception enables variable-length non-autoregressive generation by learning to both denoise and insert frames in any order.
1 Introduction
Recent advances in diffusion and flow-matching have unlocked high-fidelity image generation Esser et al. (2024); Labs (2024) and are rapidly migrating to video. Current video generation models typically adopt one of two paradigms: full-sequence generation denoises all frames jointly with full attention Wan et al. (2025); HaCohen et al. (2024); Zheng et al. (2024), while temporal autoregressive (AR) generation produces frames (or blocks of frames) sequentially in a left-to-right order Bruce et al. (2024); Song et al. (2025). Full-sequence models benefit from bidirectional attention, enabling the model to correct errors during de-noising and achieving superior generation quality. However, parallel denoising the entire sequence prevents real-time streaming, as frames cannot be returned until fully denoised. Moreover, full-sequence models use a fixed generation length and incur a quadratic attention cost in the number of frames, limiting long-term generation.
Figure 2: Flowception sampling. At each iteration, the model predicts, for each frame $i$, a velocity field and an insertion rate $\lambda_{i}$. Velocities are used to denoise frames while the insertion rates define the probability to insert a new frame to the right of existing ones. The model uses per-frame time values, set to $t_{i}\!=\!0$ when they are inserted, and reaching $t_{i}\!=\!1$ when they are fully denoised.
In contrast, AR methods enable streaming generation by making each frame immutable once generated, allowing immediate display to users while subsequent frames attend to prior outputs. However, these AR approaches suffer from critical exposure bias Zhang and Agrawala (2025); Bengio et al. (2015): training uses ground-truth frames as context, while inference conditions on the model’s own imperfect generations. This train-test mismatch prevents the model from learning to recover from its mistakes, causing minor artifacts to cascade as errors accumulate across frames, rapidly degrading video quality. Moreover, to enable KV caching, without which sampling from AR models becomes prohibitively expensive, AR methods typically use a causal attention mask which limits the expressiveness of the models. These trade-offs hinder long-term and efficient video synthesis.
In this work, we address the error accumulation and the limiting causal attention pattern of AR methods, while also avoiding the fixed-length generation requirement and reducing the computational cost of full-sequence models. To this end, we introduce Flowception , a variable-length, non-autoregressive video generation framework that interleaves two processes throughout sampling: (i) continuous flow matching denoising of existing frames, and (ii) stochastic discrete insertion of frames in between existing ones. Because the model flexibly determines the insertion locations, the same model naturally performs text-to-video, image-to-video, video-to-video, video interpolation, and scene completion by toggling which conditioning input frames are active (allow insertions to their right) or passive (no insertions). See Figure˜1 for an illustration of image-to-video and video interpolation results generated by our model.
Concretely, at each timestep the model predicts a velocity field over the existing frames and a per-frame insertion rate. By sampling based on the insertion rate, a new frame may be inserted and initialized with a sample from a unit Gaussian prior distribution. The new frame is then subsequently denoised in the context of other already instantiated and partially denoised frames. This yields a coupled ODE–jump process over variable-length sequences and supports any-order, any-length generation by design. See Figure˜2 for an illustration of the generation process.
Beyond a more flexible generative model, our interleaved insertion-denoising schedule delivers an efficiency gain, since early in sampling only a small active subset of frames is actively denoised while unrevealed frames are marginalized out. Under a linear frame insertion schedule, the computational cost of the quadratic attention term along the generation trajectory averages to a third of the attention cost incurred by a full-sequence model. Compared to AR with KV caching, Flowception shows comparable sampling cost with more robustness under low NFEs.
We conduct extensive experiments with three datasets, Tai-Chi-HD, RealEstate10K and Kinetics 600, and consider both image-to-video generation and video interpolation tasks. We find that, beyond being a more flexible framework, Flowception leads to comparable or better performance than full-sequence and autoregressive paradigms under the same compute budget, with consistent improvements in terms of FVD and VBench metrics.
Contributions:
  1. 1.
    We introduce Flowception , a theoretically grounded video generation framework, which couples frame insertions with continuous flow matching in a unified model.
  2. 2.
    We show how Flowception can solve different tasks by conditioning on any set of frames, based on their relative order alone.
  3. 3.
    We present an efficiency analysis showing an average of $3\times$ FLOPs reduction during training and ($1.5\times$ during sampling) w.r.t. full-sequence model.
  4. 4.
    We present extensive experimental results improving over full-sequence and autoregressive approaches on multiple datasets.
2 Related work
Diffusion & Flows.  Diffusion (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021; Dhariwal and Nichol, 2021) and flow models (Lipman et al., 2023; Ma et al., 2024; Esser et al., 2024) are a class of generative models that operate by learning to model the infinitesimal transitions of the probability path going from source (noise distribution) to target (data manifold). They are considered state-of-the-art generative modeling tools for modalities ranging from images (Rombach et al., 2022; Podell et al., 2024; Chen et al., 2024; Esser et al., 2024), audio (Wang et al., 2023; Levy et al., 2023), and video (Jin et al., 2025; Polyak et al., 2025). These models generate data by starting with a simple prior such as a unit Gaussian, and iteratively refining samples to recover samples from the data distribution model. The simple training recipe underlying these models proved very effective through scalability and stable training Berrada et al. (2024).
Video generation.  The success achieved by scaling text-to-image models Podell et al. (2024); Esser et al. (2024); Labs (2024) paved the way towards applying the same paradigms to video generation. Such works fall broadly into two categories. The first line of work Zheng et al. (2024); Peng et al. (2025); Wan et al. (2025) follows the full-sequence paradigm where the video is treated as one large tensor and all frames are denoised simultaneously. However, this results in prohibitively large attention computations, making the development of such methods prohibitive for long sequences.
Another line of work Deng et al. (2024); Chen et al. (2025) takes an autoregressive approach to iteratively predict the next frame (or block of frames) conditioned on previous frames. This approach is known to have difficulty generating long videos, as sampling fixes early frames once they are generated and later frames are conditioned on (imperfectly) generated earlier frames (unlike in training, where conditioning is always based on ground-truth frames), leading to error accumulation that harms long-term generation. To resolve this, some methods rely on partial noising of the context Song et al. (2025), but this prevents KV caching, making sampling prohibitively expensive. Even without KV caching, our Flowception remains significantly faster than these autoregressive methods due to its interleaved insertion and denoising mechanism, whereas autoregressive generation is fundamentally constrained by sequential frame-by-frame generation.
Another family of approaches factorizes long video generation into sparse keyframe synthesis followed by video interpolation. MovieDreamer Zhao et al. (2024) first autoregressively predicts a sequence of visual keyframes and then applies an image-to-video model to synthesize short clips around each keyframe, using CLIP features of the anchor frame rather than the last generated frame to mitigate drift across clips. ART-V Weng et al. (2024) instead performs frame-wise autoregressive generation in a text-image-to-video (TI2V) setting: each new frame is produced by a diffusion model conditioned on both the input keyframe and the previously generated frames, with a masked diffusion mechanism that selectively reuses information from the given frame to reduce appearance drift over long horizons.
Several other directions have been explored to reduce complexity. These include the use of highly compressive tokenizers HaCohen et al. (2024); NVIDIA et al. (2025), hierarchical or packed representations that compress remote frames before attending to them Jin et al. (2025); Zhang and Agrawala (2025), and architectural changes that distribute or route attention over multiple context sources Fan et al. (2025); Cai et al. (2025); Song et al. (2025); Zhao et al. (2024). However, these approaches still depend on design choices about which context sources or time ranges to emphasize, rather than fully learning relevance end-to-end.
To enable multiple use cases, such as image-to-video and video-to-video, some works Wan et al. (2025); HaCohen et al. (2024) concatenate a frame-wise mask to the input of the model, which allows to condition on clean frames at various positions in the video. Similarly, Flowception is capable of performing various tasks by conditioning on multiple input frames, and optionally constraining insertions to be to the right of some of those.
3 Flowception
We propose a natural combination of the continuous Flow Matching (Lipman et al., 2023) and the discrete Edit Flow (Havasi et al., 2025) frameworks, with the goal of flexible video generation where existing frames are simultaneously denoised and new frames are inserted during the generation process. This is our Flowception model. We first describe the parameterization and the generation procedure for this model, then derive a training scheme that adheres to the generation procedure. We will present notation loosely with high-level intuitive explanations; full derivations are in the supplementary material.
3.1 Setup and model parameterization
Flowception operates in the space of variable-length sequences of frames $\mathcal{X}=\bigcup_{n=0}^{\infty}\mathbb{R}^{n\times H\times W\times C}$, where $n$ denotes the length of each sequence, and $H,W,C$ are the fixed height, width, and channel dimensions for each frame. We denote $\ell(X)$ as the length of a sequence $X\in\mathcal{X}$.
During generation, Flowception transports sequences and time values $(X,t)$ from an initial noise distribution $p_{\text{src}}$ to a data distribution $p_{\text{data}}$ through a continuous-time framework. We handle variable length by assigning per-frame time values, collected in a vector $t\in\mathcal{T}=\bigcup_{n=0}^{\infty}[0,1]^{n}$, associating each frame to a different noise level. Hence the input of a Flowception model is a sequence of frames $X\in\mathcal{X}$ and a vector of per-frame time values $t\in\mathcal{T}$, where $X$ and $t$ are the same length.
To incorporate sequence length changes, we make use of an insertion operation (Havasi et al., 2025). Given a sequence $X\in\mathcal{X}$, an insertion index $i$, and a noise frame $\varepsilon\sim\mathcal{N}(0,I)$, we define
${{\text{ins}}}(X,i,\varepsilon)=(X^{1},\dots,X^{i},\varepsilon,X^{i+1},\dots,X^{n}),$ (1)
where the $X^{i}\in\mathbb{R}^{H\times W\times C}$ are frames in the sequence. The insertion operation inserts a 0-SNR frame which will then be denoised following the flow matching framework.
Frame insertions and flows form the primitive operations that we use during generation. Hence the output of a Flowception model with parameters $\theta$ has two components at every position $i\in\{1,\dots,\ell(X)\}$:
  1. 1.
    An insertion rate $\lambda_{i}^{\theta}(X,t)\in\mathbb{R}_{\geq 0}$, which predicts the number of missing frames.
  2. 2.
    A velocity $v_{i}^{\theta}(X,t)\in\mathbb{R}^{H\times W\times C}$, for denoising the existing frames.
These two components allow the model to simultaneously denoise an existing sequence of frames and interpolate or extend the existing frame sequence.
3.2 Generation procedure
We start the generation process by sampling a fixed number of starting frames $n_{\text{start}}$,
$p_{\text{src}}(X)=\textstyle\prod_{i=1}^{n_{\text{start}}}\mathcal{N}(X^{i};0,I),$ (2)
and assigning each frame the time $t^{i}=0$. We use a non-zero number of starting frames to start the generation process, otherwise the initial generation steps will simply be inserting noise frames anyway.
We then iteratively apply a transport step until all frames are clean, i.e., all $t^{i}=1$. Each transport step simultaneously flows the existing frames while also potentially inserting new frames with probability given by the insertion rates.
Global time value.  The stopping criterion for the flow is when $t^{i}=1$ for all frames, i.e. we no longer modify any frame. To impose a stopping criterion for the insertions, we introduce a global time value $t_{g}$ which starts at $t_{g}=0$ and frame insertions are only allowed while $t_{g}<1$. Implementation-wise, we don’t feed $t_{g}$ explicitly into the model, and is only tracked for sampling. This closely follows the design choice of OneFlow Nguyen et al. (2025) where insertion rates are not conditioned on time. Accordingly, with a constant step size during inference, the maximum number of iterations is twice the amount needed for $t_{g}$ to arrive to $1$.
Scheduler.  We impose a distribution on the fraction of visible frames based on a monotonic scheduler $\kappa(t_{g})$ that is a function of the global time value, with $\kappa(0)=0$, $\kappa(1)=1$. Given a data sequence $X_{1}\sim p_{\text{data}}$, the probability of each non-starting frame being in the sequence at a global time $t_{g}$ is, for each frame $X^{i}\in X_{1}$,
$\mathbb{P}(X^{i}\text{ in }X\;|\;\text{ global time }t_{g})=\kappa(t_{g}).$ (3)
We will primarily work with the linear scheduler $\kappa(t)=t$.
Transport step.  Given a sequence $X$ with per-frame times $t$, each transport step performs the following two operation simulatenuosly at all positions $i\in\{1,\dots,\ell(X)\}$:
  1. 1.
    Apply the flow step
    $X^{i}=X^{i}+hv_{i}^{\theta}(X,t).$ (4)
  2. 2.
    With probability $h\tfrac{\dot{\kappa}(t_{g})}{1-\kappa(t_{g})}\lambda_{i}^{\theta}$, insert a new frame
    $X={{\text{ins}}}(X,i,\varepsilon),\quad t={{\text{ins}}}(t,i,0).$ (5)
The ratio $\tfrac{\dot{\kappa}(t_{g})}{1-\kappa(t_{g})}$ ensures insertions occur in aligment with the distribution of visible frames imposed by the scheduler.
To summarize, we start with a fixed number of $n_{\textrm{start}}$ starting frames initialized as noise, advance a global time value $t_{g}$ from 0 to $1$, and perform frame insertions along the way following the scheduler $\kappa(\cdot)$ in Equation˜3. Noise frames are inserted and then subsequently denoised, naturally producing per-frame time values $t_{i}$ and with a delay in the time values of later-inserted frames (i.e., $t_{i}\leq t_{g}$). We stop the generation when all frames have reached $t_{i}=1$. An illustration of a transport step is shown in Figure˜2. A pseudocode implementation is in the supplementary material.
3.3 Training procedure
The Flowception generation procedure induces a particular distribution over the visible frames and their time values. Frames may be at different time values, and some may not be inserted yet. We must align with this distribution of time values and missing frames during training so that there is no distribution mismatch between training and generation. Despite this complexity, the distribution is controlled completely by the scheduler $\kappa$. In this section, we describe a simple training recipe to easily sample all time values and the states of each frame during training.
Extended time values.  Extended time values are denoted $\tau$ and can take values outside of the interval $[0,1]$, so we define a clip operation that results in real time values,
$t=\text{clip}(\tau):=\max\{0,\min\{1,\tau\}\}.$ (6)
Firstly, after the global time value reaches $t_{g}=1$, while no insertion of new frames can occur, existing frames may still need more steps to be fully denoised. Thus we introduce an extended global time $\tau_{g}\in[0,2]$, with $t_{g}=\text{clip}(\tau_{g})$.
Secondly, while $\tau_{g}\leq 1$, insertions occur according to the scheduler $\kappa(\cdot)$. For the linear scheduler—for the general case, please see the detailed derivations in the supplementary material—new frames are inserted uniformly across the time interval $t_{g}\in[0,1]$. Therefore, the time delay between the global extended time and per-frame extended time values follows:
[…]
where $t_{i}=\text{clip}(\tau_{i})$ computes the real per-frame times. Note that $\tau_{i}\in[-1,2]$ according to (7), When $\tau_{i}<0$, the frame is yet to be inserted and thus in a “deleted” state.
Figure˜3 illustrates the different frame states according to the extended time schedule. Each frame can either
  • be in deleted state ($\tau_{i}<0$),
  • exist and be in flow state ($0\leq\tau_{i}<1$) , or
  • be frozen and in terminal state ($\tau_{i}\geq 1$).
During training, we sample noisy sequences using
$\tau_{g}\sim p(\tau_{g}),\quad\tau_{i}=\tau_{g}-u_{i},$ (8)
$t_{i}=\text{clip}(\tau_{i}),\quad X_{1}^{\text{visible}}=(X^{i}\in X_{1}|\tau_{i}>0),$ (9)
$X=tX_{1}^{\text{visible}}+(1-t)X_{0},\quad X_{0}\sim\mathcal{N}(0,I),$ (10)
where $X_{1}\sim p_{\text{data}}$ is a target sequence sample.
Figure 3: Illustration of the extended time scheduler for Flowception training. In Flowception , each frame has its own denoising time which depends on its insertion time. The global extended time $\tau_{g}$ progresses from 0 to $2$, where insertion of new frames only occur when $\tau_{g}<1$. Starting frames (in blue) are instantiated at $\tau_{g}=0$, other frames (in orange) are inserted later (when $\tau_{g}>0$) and thus have a delay. With a linear scheduler, the insertion delays follow a uniform distribution.
Insertion loss.  At each position $i$ of the noisy sequence $X$, we denote $k^{i}$ as the number of missing frames to the right, between positions $i$ and $i+1$. Following OneFlow Nguyen et al. (2025), the loss for training the insertion output $\lambda_{i}$, which predicts the number of missing frames, is given by the negative log-likelihood of the Poisson distribution
$\mathcal{L}_{\mathrm{ins}}=\sum_{i=1}^{\ell(X)}\lambda_{i}^{\theta}(X,t)\;-\;k^{i}\,\log\lambda_{i}^{\theta}(X,t).$ (11)
Velocity loss.  At each position $i$ of the noisy sequence $X$, we learn to denoise the frame with the standard Flow Matching loss (Lipman et al., 2023),
$\mathcal{L}_{\text{vel}}=\left\lVert v^{\theta}(X,t)-(X_{1}^{\text{visible}}-X_{0})\right\rVert^{2}.$ (12)
3.4 Practical considerations & implications
Architecture with per-frame time conditioning.  Differently from other frameworks, our model has two prediction heads: the dense velocity prediction head and the insertion rate prediction head. In order to enable different timesteps per frame, we change the AdaLN Peebles and Xie (2023) modulation to operate on a per-frame basis. To distinguish between noisy and input/conditioning frames, we inflate the channel dimension of the inputs so that the first $C$ channels contain the noisy frames while the second group contains the input/conditioning frames (or zero padding). Finally, for the rate prediction head, a learnable token is concatenated to each frame’s tokens before being projected with a simple MLP and an exponential activation.
Figure 4: Flowception natively supports different tasks. By choosing active and passive context frames the model can be used for image-to-video, video-to-video, text-to-image, frame interpolation, text-to-video, and scene completion.
Supporting different tasks.  Since the frames present at time $t$ are only specified by their relative order, Flowception supports multiple tasks by feeding different context frames. We distinguish two sorts of context frames: active ones which can induce insertions, and passive ones which do not. As illustrated in Figure˜4, mixing these context frames in different ways allows for image-to-video, video-to-video interpolation and scene completion without needing to specify the size of the gaps but only the relative order of the context frames.
Computational cost.  Early in sampling, only a small subset of frames is visible, which reduces the computational cost of these denoising steps compared to full-sequence denoising. Under linear $\kappa$ we have $\mathbb{E}_{p(\tau_{g})}[\kappa(\tau_{g})^{2}]=\int_{0}^{1}\tau_{g}^{2}d\tau_{g}=1/3$, and so the expected attention FLOPs are $1/3$ of full-sequence Flow Matching, assuming an equal number of denoising steps used in both models. During sampling, as the latest flow update occurs at $\tau_{g}=2$, our model uses at most twice the flow iterations of the full-sequence model assuming the same time discretization, i.e. $2/3$ the FLOP count of full-sequence Flow Matching. Compared to AR models, Flowception has a comparable sampling cost with more robustness under low NFEs. A more detailed analysis is available in the supplementary material.
4 Experiments
Table 1: Image-to-Video generation results. Models trained and sampled at 256 resolution. Comparing our Flowception approach to the Full Sequence and (causal) Autoregressive baselines. VBench Quality metrics include imaging quality, background consistency, aesthetic quality, motion smoothness, subject consistency, and dynamic degree. Best results for each dataset are highlighted.
VBench Metrics Quality
[…]
Kinetics-600
Full Seq 37.09 94.75 39.42 99.39 92.46 44.35 204.65
Autoreg. 38.77 92.69 38.17 98.11 85.82 54.66 201.34
Flowception 41.92 96.96 42.05 99.21 94.74 47.07 164.73
Tai-Chi-HD
Full Seq. 47.48 94.42 54.93 99.26 92.18 18.61 27.30
Autoreg. 47.15 95.93 54.98 99.37 93.41 21.23 25.30
Flowception 48.42 95.93 54.96 99.67 94.43 20.02 25.21
RealEstate10K
Full Seq. 50.11 93.48 44.53 99.08 85.85 81.64 26.17
Autoreg. 48.55 93.84 44.48 99.16 87.29 72.60 47.48
Flowception 51.18 96.93 48.09 99.30 87.02 78.59 21.80
Figure 5: Image-to-video generation using model trained on Tai-Chi-HD. We show two examples (one per row) of generated video frames along the insertion time of each frame. In general, frames inserted early, with $t\approx 0$, define the movement dynamics (large changes w.r.t. the context frame), while later frames smoothly interpolate motion, resulting in smaller changes w.r.t. to neighboring frames.
4.1 Experimental setup
Model architecture.  Our architecture builds on the popular DiT diffusion transformer architecture Peebles and Xie (2023), with minimal changes to enable insertion rate predictions. We equip each attention block with an additional learnable token which is replicated for each frame and takes part in the attention computation, ultimately each of these tokens maps to the non-negative rate prediction using an exponential activation. Our baseline model is made of 38 attention blocks with a hidden dimension of 1536, each attention head has a dimension of 64, resulting in 24 total heads, for a model with approximately 2.1B learnable parameters. Positional embeddings follow the VideoROPE method Wei et al. (2025). We enable full bidirectional attention between frames and optionally concatenated text tokens in an MMDiT Esser et al. (2024) fashion.
Datasets.  We evaluate our model on three different datasets. We used two narrow-domain datasets: Tai-Chi-HD Siarohin et al. (2019) and RealEstate10K Zhou et al. (2018), as well as the class-structured Kinetics 600 dataset Carreira et al. (2018). Unless specified otherwise, we train our models at $256$ resolution and generate up to 145 frames at 16 FPS. Furthermore, we trained a general domain model using a proprietary dataset of around 20M videos ranging approximately from 20 to 240 seconds. We show results for this model in the supplementary material. We use the LTX autoencoder HaCohen et al. (2024), which has a spatial downsampling factor of 32 and temporal downsampling factor of 8. To correctly support variable length video decoding, we modify the decoder architecture to propagate a frame validity mask which is used to mask out invalid frames in the hidden layers, preventing border effects where padding frames leak into the video. When prompting models with one or more video frames to generate a video, we choose these reference frames from the validation set.
Metrics.  To evaluate the performance of our models, we rely on the standard FVD metric Unterthiner et al. (2019), computed with respect to a subset of 5k videos from the training set, as well as selected metrics from VBench Huang et al. (2024). Specifically, we report imaging and aesthetic quality, background and subject consistency, motion smoothness and dynamic degree.
Baselines.  We compare our method with popular video generation paradigms including (i) full-sequence generation where all frames are flowed simultaneously, sharing a global timestep and bidirectional attention over all frames, and (ii) per-frame autoregressive generation where the model iteratively predicts the next frame(s) conditioned on the previously available ones. For autoregressive generation, we experiment with both bidirectional and causal attention (which allows KV caching for efficient inference).
We sample videos from Flowception and baselines without using guidance, unless specified otherwise. The complete videos from which we show selected frames in Figures˜1, 5 and 6 are included in the supplementary material to better appreciate the video quality.
4.2 Main results
Image-to-Video (I2V) generation.  In Table˜1 we compare our Flowception approach to the full-sequence and autoregressive baselines for the I2V task where we condition on the first frame to generate the rest of the video. We set the generated video length to 145 frames. Overall, we find the autoregressive and full-sequence model to give fairly similar results. For the VBench metrics, Flowception leads to better results in most cases, or second best otherwise. For video quality as measured in FVD, Flowception improves results across the board.
In Figure˜5 we show two qualitative examples of I2V generation, which demonstrate the ability of the model to generate smooth video sequences with substantial motion and appearance consistency. Additionally, we observe an emergent “coarse-to-fine” structure in the order of generated frames: early on in the generation process far-away frames are generated that tend to vary and define the overall motion of the video, while later frames mostly interpolate between them.
Figure 6: Video interpolation results obtained with with Flowception trained on the Kinetics-600 dataset. Context frames are highlighted with dashed lines, insertion times of other frames are marked using a color map and printed on each frame.
Video interpolation.  Another interesting task directly supported by Flowception is video interpolation, where rather than a single starting frame, we provide an ordered set of context frames as conditioning, and let the model interpolate the video freely between these. Differently from other frameworks Wan et al. (2025); HaCohen et al. (2024), Flowception does not need to be given the number of frames between successive conditioning frames, but it rather inserts frames as it sees fit, resulting in a more flexible interpolation capability. We illustrate this use-case in Figure˜6. We observe that the model is able to adapt the number of insertions to generate coherent scenes with smooth transitions. For instance, in the second row, no frame is inserted after the penultimate context frame in order to preserve motion continuity.
Table 2: Comparing insertion schemes. Using the Flowception rate prediction and baselines that insert frames in a random order, in a hierarchical scheme (see text for detail), and left-to-right. Models trained on RealEstate10K.
Insertion method FVD Motion Dynamic
Random 25.03 99.09 70.68
Hierarchical 23.94 99.10 71.20
Left-to-right 23.61 99.03 73.04
Flowception rate prediction 21.80 99.30 78.59
4.3 Ablation experiments
Impact of learned insertions.  To assess the importance of the insertion head that learns to predict where to insert frames, we compare it to inserting frames in a data-independent manner. In particular, we first determine the video length $n$. We then consider inserting frames in a random order, using a “hierarchical” scheme where we iteratively insert them in the middle of the largest interval in $\{1,\dots,n\}$ where no frame has been inserted yet, or using a left-to-right pattern as used in AR models. We present the results of this experiment in Table˜2. We find that across metrics, results improve as we move from the random pattern to the hierarchical pattern, to the left-to-right, and finally to the data-driven insertion patterns in Flowception .
Note that for the left-to-right pattern, the FVD results in Table˜2 (23.61) are much better than the FVD for the AR baselines in Table˜4 (47.48 and 45.13). We hypothesize that this difference is due to the fact that the AR baselines denoise the next frame once previous ones are already fully denoised and committed to, while in Flowception new frames start denoising while previous ones are still not fully denoised (more similar to the full-sequence baseline which yields a FVD of 26.17 on RealEstate10K).
Insertion guidance.  Classifier-free guidance (CFG) is commonly used to achieve better prompt alignment and image quality in diffusion and flow models Ho and Salimans (2021). Similarly, in Flowception , we can perform CFG on the insertion rates. Following CFG, we randomly drop the conditional information $c$ during training, and define the guided update as
$\lambda^{\text{cfg}}(X_{t}|c)=\lambda(X_{t}|c)^{w_{s}}\lambda(X_{t})^{1-{w_{s}}},$ (13)
where $w_{s}\geq 1$ is the guidance scale for the insertion rate. Figure˜7 presents a scatter plot comparing video lengths for I2V obtained with and without guidance, initiating generation from the same seed for each starting frame. For efficiency we limit the number of frames to 20 in this experiment. The result clearly indicates that using rate guidance ($w_{s}\!=\!5$) biases the model towards generating longer videos.
Figure 7: Impact of rate guidance on I2V video length. Samples obtained without guidance ($w_{s}\!=\!1$, horizontal) and with guidance ($w_{s}\!=\!5$, vertical) on RealEstate10K, using the same seed for each conditioning frame.
Without guidance, or using lower values, we find insertion guidance to prevent problems of under-insertions which can result in a choppy transition between two frames. This is confirmed in Table˜3 where motion smoothness increases while dynamic degree decreases as $w_{s}$ increases.
Table 3: Effect of rate guidance on motion smoothness. We report dynamic degree and motion smoothness for various values of the insertion rate guidance parameter $w_{s}$.
$w_{s}$ FVD Motion Dynamic
1.0 21.80 99.30 78.59
2.0 22.69 99.31 78.61
5.0 25.30 99.33 77.78
Causality in AR models.  In the literature, most autoregressive video generation models make use of frame-wise causal attention Deng et al. (2024); Chen et al. (2025), which allows for KV caching and therefore efficient inference for autoregressive generation. Other works improve on quality and long term rollout consistency by using bidirectional attention in the context and applying frame-wise noising for each frame Song et al. (2025). Such an approach, while effective, significantly raises inference time expected compute from $O(L^{2}n^{2})$ to $O(L^{2}n^{3})$ where $n$ is the number of video frames, and $L$ is the number of tokens per frame. A more detailed analysis of FLOPs is provided in the supplementary material.
We provide a comparison between both autoregressive methods in Table˜4. The results demonstrate slightly better performance of the non-causal attention pattern, outperforming the causal one on every metric. Flowception yields further improvements over the non-causal AR baseline on all metrics but subject consistency.
Table 4: Comparison of autoregressive attention patterns. Autoregressive baseline with differnt attention amskes and comparison to Flowception in terms of VBench metrics and FVD, Models trained on Real Estate10K.
AR causal AR non-causal Flowception
FVD 47.48 45.13 21.80
Imaging quality 48.55 48.70 51.18
Background consistency 93.84 93.88 96.93
Aesthetic quality 44.48 45.28 48.09
Motion smoothness 99.16 99.20 99.30
Subject consistency 87.29 87.46 87.02
Dynamic degree 72.60 73.52 78.59
Figure 8: Comparing local attention variants. We report FVD on RealEstate10K after training for 300k iterations with different attention context windows. For each frame, its context is made of itself and the previous/next context window frames.
Local vs. global attention.  Flowception generates a video by progressively inserting new frames. Empirically, we observe that Flowception tends to insert distant frames first, and then progressively “fills in the blanks”. This suggests that Flowception could be more amenable to using efficient local attention patterns, as far-away frames can still attend to each other early in the denoising process when the sequence is still short. To assess whether Flowception is more amenable to using local attention windows than the full-sequence baseline, we compare models using global attention and using local attention windows restricted to the previous and next $K$ frames in the sequence. The results in Figure˜8 indicate that Flowception suffers far less from using local attention windows than the full-sequence model, presumably due the communication between distant frames early in the Flowception denoising process, when the intermediate frames have not yet been inserted, even when using small attention windows.
5 Conclusion
We presented Flowception , a novel video generation model that interleaves denoising existing frames with inserting new ones at appropriate locations, enabling efficient variable-length non-autoregressive generation. Our method outperforms full sequence and autoregressive baselines in both generation quality and training efficiency. Flowception naturally supports various prediction tasks including image-to-video, video-to-video generation, and frame interpolation. Our work offers a promising alternative to standard autoregressive and full-sequence approaches for long-term generation and flexible editing. While training on partial sequences allows for these emergent behaviors,it doubles the number of iterations in order to ensure complete denoising of all frames. Exploring different interleaved schedules for improved efficiency is an exciting future direction.